OpenStudy (anonymous):

Substitute identities on the left to match the right hand side of the equation. cot x sec4x = cot x + 2 tan x + tan3x

jimthompson5910 (jim_thompson5910):

so the equation is \[\large \cot(x)\sec(4x) = \cot(x) + 2\tan(x) + \tan(3x)\] right?

OpenStudy (anonymous):

ugh not quite let me use the wonderful tools this site provide to make it easier

OpenStudy (anonymous):

cot(x) sec^4x = cot(x) + 2tan (x) + tan^3x

OpenStudy (anonymous):

except that I couldn't figure out how to use the tools so I did that instead

jimthompson5910 (jim_thompson5910):

oh ok, so it's \[\large \cot(x)\sec^4(x) = \cot(x) + 2\tan(x) + \tan^3(x)\]

jimthompson5910 (jim_thompson5910):

one sec while I think on how to do this one

OpenStudy (anonymous):

here does this help? file:///C:/Users/IBMT61/Pictures/05.03%20Proving%20Trigonometric%20Equations.htm

OpenStudy (anonymous):

sure take your time

jimthompson5910 (jim_thompson5910):

no I can't access that link at all (it's local to your computer)

jimthompson5910 (jim_thompson5910):

it's fine though, I can figure it out

OpenStudy (anonymous):

hmm let me try this then...

OpenStudy (anonymous):

am I a genius in a small form? :D

jimthompson5910 (jim_thompson5910):

oh that works much better, thanks

OpenStudy (anonymous):

The English nerd doesn't break the tech department for once...praise the gods

OpenStudy (anonymous):

Isn't it worded really strangely though? If it wasn't for the fact that I KNOW I have to chance the LEFT side I'd be confused and having to wait for my teacher to ask that

OpenStudy (anonymous):

change*

jimthompson5910 (jim_thompson5910):

well the instructions are worded a bit oddly

jimthompson5910 (jim_thompson5910):

but normally you start with the more complicated side and you simplify it with identities to change it to the other side you didn't pick

OpenStudy (anonymous):

This is all that the lesson has...everything else is in a generic online college text book...my high school lessons page itself however only offers THIS

jimthompson5910 (jim_thompson5910):

Here's how I did it. I started on the right side (because it was much more complicated) and simplified it to get the left side. I did NOT alter the left side at all.

jimthompson5910 (jim_thompson5910):

\[\large \cot(x)\sec^4(x) = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \cot(x)\sec^4(x) = \frac{1}{\tan(x)} + 2\tan(x) + \tan^3(x)\] \[\large \cot(x)\sec^4(x) = \frac{1}{\tan(x)} + 2\tan(x)*\frac{\tan(x)}{\tan(x)} + \tan^3(x)*\frac{\tan(x)}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{1}{\tan(x)} + \frac{2\tan^2(x)}{\tan(x)} + \frac{\tan^4(x)}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{1+2\tan^2(x)+\tan^4(x)}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{(1+\tan^2(x))^2}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{(\sec^2(x))^2}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{\sec^4(x)}{\tan(x)}\] \[\large \cot(x)\sec^4(x) = \frac{1}{\tan(x)}*\sec^4(x)\] \[\large \cot(x)\sec^4(x) = \cot(x)\sec^4(x)\]

OpenStudy (anonymous):

The thing is that my course specifically states that I have to change the left side...as you can see in that png file I sent you last...I'm sure that means they want me to do things in a much harder way...of bloody course they do

jimthompson5910 (jim_thompson5910):

hmm, there's probably a way to do that, let me think

jimthompson5910 (jim_thompson5910):

oh duh, you just follow my work backwards lol

jimthompson5910 (jim_thompson5910):

here let me see if I can't translate what I mean

OpenStudy (anonymous):

Merci beaucoup

jimthompson5910 (jim_thompson5910):

\[\large \cot(x)\sec^4(x) = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{1}{\tan(x)}*\sec^4(x) = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{\sec^4(x)}{\tan(x)} = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{(\sec^2(x))^2}{\tan(x)} = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{(1+\tan^2(x))^2}{\tan(x)} = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{1+2\tan^2(x)+\tan^4(x)}{\tan(x)} = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{1}{\tan(x)} + \frac{2\tan^2(x)}{\tan(x)} + \frac{\tan^4(x)}{\tan(x)} = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \frac{1}{\tan(x)} + 2\tan(x) + \tan^3(x) = \cot(x) + 2\tan(x) + \tan^3(x)\] \[\large \cot(x) + 2\tan(x) + \tan^3(x) = \cot(x) + 2\tan(x) + \tan^3(x)\] The identity is verified.

jimthompson5910 (jim_thompson5910):

oh btw that "duh" was just me thinking out loud and me thinking "oh yeah I should have known that", it wasn't aimed at you (I'm sure you know that though) anyways, i hope that all makes sense

OpenStudy (anonymous):

I gotta read through that and see if I get to where your mind is...but I'll let you know if I feel stuck at all...and thank you for this! I wish I could give you a medal but I don't know how :(

jimthompson5910 (jim_thompson5910):

you just click "best response" but I don't mind either way

jimthompson5910 (jim_thompson5910):

and yeah, feel free to ask any questions you have about any of that above